# On computational complexity and average-case hardness of shallow-depth Boson Sampling

### v2

@@ Appendix E Proof of Theorem 4 (average-case hardness)
For the polynomial interpolation, we employ the Robust Berlekamp-Welch algorithm that requires the NP oracle, recently proposed in Ref. [28].
Theorem 10 (Robust Berlekamp-Welch [28]).
Let $P(x)$ be a degree $d$ polynomial in $x$. Suppose there is a set of points $D=\{(x_{i},y_{i})\}$ such that $|D|=O(d^{2})$ and $\{x_{i}\}$ is equally spaced in the interval $[0,\Delta]$. Suppose also that each points $(x_{i},y_{i})$ satisfies
@@
Appendix F Proof of Lemma 2
Let $\bar{p}_{\bm{s}}(C)$ be the output probability distribution from the approximate sampler $\mathcal{S}$ with the given linear optical circuit $C$. Also, let $\mathcal{C}_{M,N}$ be the set of collision-free outcomes of Boson Sampling, for mode number $M$ and photon number $N$. Then $\bar{p}_{\bm{s}}(C)$ satisfies
@@ Appendix F Proof of Lemma 2
for all $k>2$. Also, using Stockmeyer’s algorithm [47] whose complexity is in $\rm{BPP}^{\rm{NP}}$, obtaining the estimate $\tilde{p}_{\bm{s}}(C)$ of $\bar{p}_{\bm{s}}(C)$ satisfying
@@
in polynomial time in $N$ and $\alpha^{-1}$ is in $\rm{BPP}^{\rm{NP}^{\mathcal{S}}}$. Using $\operatorname*{\mathbb{E}}_{\bm{s}\sim\mathcal{G}_{M,N}}[\bar{p}_{\bm{s}}(C)]=\binom{M}{N}^{-1}\sum_{\bm{s}\in\mathcal{C}_{M,N}}\bar{p}_{\bm{s}}(C)\leq\binom{M}{N}^{-1}$,
@@
Appendix G On average-case hardness including collision outcomes
@@
As we have remarked in the main text, extending our hardness result of shallow-depth Boson Sampling to the saturated regime $M=\Theta(N^{\gamma})$ for $21\leq\gamma<$ 2 remains a crucial open problem, and to do so, one also needs to consider collision outcomes in the average-case hardness argument. In this appendix, we present a possible approach that can be further improved to the average-case hardness result including collision outcomes.
@@
Our intuition is that while most of the outcomes in the saturated regime are not collision-free, the key point is that most outcomes are expected to be “nearly” collision-free, in the sense that for most of the outcomes, a large portion of the measured modes are occupied only by a single photon. Also, we can establish worst-to-average-case reduction across different system sizes (i.e., mode number & photon number), i.e., the system size for the worst-case hardness result can be smaller at most polynomially than the system size for the average-case hardness result. Hence, by post-selecting outcomes with the number of measured single photons at least as large as the photon number required in the worst-case problem (in Theorem 3, which requires a collision-free outcome), one would possibly establish worst-to-average-case reduction from the randomly chosen outcome over all possible outcomes.
@@
More specifically, let the mode number $M=\Theta(N^{\gamma})$ for any $\gamma\geq 1$. Let us define $\,\mathcal{Q}_{M,N}$ as the uniform distribution over all the possible outcomes of Boson Sampling with $M$ modes and $N$ photons, such that each outcome $\bm{s}\sim\mathcal{Q}_{M,N}$ is an $M$-dimensional output configuration vector $\bm{s}=(s_{1},\dots,s_{M})$ with the constraint $\sum_{i=1}^{M}s_{i}=N$. Also, we denote the number of $s_{i}=1$ in $\bm{s}$ as the number of singletons throughout this appendix. Here, let us assume that there exists $N_{0}=cN^{\lambda}$ for a constant $c$ and $10<\lambda<$ 1 such that it is likely that the number of singletons is more than $N_{0}$; we later provide the intuition to justify this assumption at the end of this appendix.
@@
Hereafter, we denote $(M_{0},N_{0})$ and $(M,N)$ as the system sizes (mode number & photon number) for the worst-case hardness (i.e., Theorem 3) and average-case hardness (we aim to prove), respectively. More specifically, given a worst-case circuit $C_{0}$ in $M_{0}=\Omega(N_{0})$ mode circuit architecture $\mathcal{B}\mathcal{B}^{*}$ and a fixed $M_{0}$-mode and $N_{0}$-photon collision-free outcome $\bm{s}_{0}$ depicted in Theorem 3, by our assumption, one can find a photon number $N$ that satisfies $N_{0}=cN^{\lambda}$ for a constant $c$ and $10<\lambda<$ 1, and mode number $M=\Theta(N^{\gamma})$ for $\gamma\geq 1$, such that most of the outcomes over $\mathcal{Q}_{M,N}$ have at least $N_{0}$ singletons. In the following, we show a possible approach for worst-to-average-case reduction for general outcomes, by describing how one can estimate $p_{\bm{s}_{0}}(C_{0})$ to within desired imprecision given oracle access to the well-estimated values $p_{\bm{s}}(U)$ for average-case outcomes $\bm{s}\sim\mathcal{Q}_{M,N}$ and $M$-mode average-case circuits over $U\sim\mathcal{H}_{\mathcal{A}}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$.
@@
We now argue how to estimate the worst-case value $p_{\bm{s}}(C)$ and the average-case value $p_{\bm{s}_{1}}(U_{0})$. First, estimating the average-case value $p_{\bm{s}_{1}}(U_{0})$ can be done by similarly accessing the oracle, but now we input the outcome $\bm{s}_{1}$ and the circuit $U_{0}$ instead. Hence, the success probability for the oracle would increase at most the total variation distance between the distribution of $\bm{s}_{1}$ and $\mathcal{Q}_{M/2,N-N_{0}}$, and if this can be sufficiently bounded, we can obtain the estimated value of $p_{\bm{s}_{1}}(U_{0})$ with high probability. Also, one can estimate the value $p_{\bm{s}}(C)$ by the polynomial interpolation, as in the proof of Theorem 4. Specifically, by parameterizing the circuit $U$ as $U(\theta)$, where each gate of $U$ (except those in $U_{0}$) is perturbed via the Cayley transform using the corresponding gate from $C$, one can similarly infer the value $p_{\bm{s}}(C)$ through oracle access to $p_{\bm{s}}(U(\theta))$ and applying polynomial interpolation (see Appendix E for details).
@@
We highlight the necessary ingredients that need to be addressed for the successful worst-to-average-case reduction by the above process. First, to make sure that we can also well-estimate $p_{\bm{s}_{1}}(U_{0})$ with high probability over $\bm{s}_{1}$ via the oracle access, one needs to justify that the distribution of $\bm{s}_{1}$ is sufficiently close to the distribution $\mathcal{Q}_{M/2,N-N_{0}}$. To help understand how the distribution of $\bm{s}_{1}$ can be close to $\mathcal{Q}_{M/2,N-N_{0}}$, consider the limiting case that $N_{0}=0$. Here, given that $M$ is much larger than $N$ (but still can be $M=\Theta(N)$), sampling $\bm{s}_{1}$ is equivalent to sampling from $\mathcal{Q}_{M,N}$ and post-selecting outcomes that have all the non-zero elements in the first $M/2$ mode, which yields the uniform distribution $\mathcal{Q}_{M/2,N}$. Therefore, given that $M$ is much larger than $N$ and $N_{0}$ is sufficiently small relative to $N$, we expect that the distribution of $\bm{s}_{1}$ would not largely deviate from the distribution $\mathcal{Q}_{M/2,N-N_{0}}$.
@@
Also, as the final imprecision level increases by the estimated value of $p_{\bm{s}_{1}}(U_{0})^{-1}$, it must be promised that $p_{\bm{s}_{1}}(U_{0})$ is lower bounded. More precisely, $p_{\bm{s}_{1}}(U_{0})$ must be at least as large as the allowed imprecision level of average-case estimation (i.e., imprecision level of the oracle $\mathcal{O}$), commonly referred to as the anti-concentration property [1, 40, 56, 57, 58, 59]. Nevertheless, we believe that $p_{\bm{s}_{1}}(U_{0})$ is at least as large as the allowed imprecision level for most of $U_{0}$, based on two reasons: First, the imprecision level required by our current average-case hardness result is sufficiently small. Also, although this has only been demonstrated in finite-sized numerical experiments, random circuits for our circuit architecture closely resemble global Haar-random unitary circuits [46], which are strongly conjectured to exhibit the anti-concentration property [1].
@@
We conclude this appendix by presenting an example that would offer an insight that, for a large portion of outcomes of Boson Sampling, a sufficient number of singletons would exist. In the following, we present the classical “balls-into-bins” problem where we throw $N$ balls into $M$ bins uniformly at random, which is indeed the classical analogue of Boson Sampling. We show that, given that $M=\Theta(N^{\gamma})$ with $\gamma\geq 1$, it is highly unlikely that the number of bins getting exactly one ball is less than $cN^{\lambda}$ for a constant $c$ and $10<\lambda<$ 1.
@@
Let us first denote the random variable $X_{i}$ as the number of balls in $i$-th bin (such that $\sum_{i=1}^{M}X_{i}=N$), when we throw $N$ balls into $M$ bins uniformly at random. Here, our goal is to verify that over the experiments, it is highly unlikely to observe the outcomes that the number of $X_{i}=1$ is smaller than $cN^{\lambda}$. Since all the variables $X_{i}$ are correlated due to the constraint $\sum_{i=1}^{M}X_{i}=N$, instead, we consider i.i.d. random variable $Y_{i}$ each drawn from Poisson distribution $P(\frac{N}{M})$, such that $\operatorname*{\mathbb{E}}[\sum_{i=1}^{M}Y_{i}]=N$ and $\Pr[\sum_{i=1}^{M}Y_{i}=N]=e^{-N}\frac{N^{N}}{N!}=\Theta(\frac{1}{\sqrt{N}})$. The motivation for using Poissonian random variables is that the distribution of $(Y_{1},\dots,Y_{M})$ conditioned on the sum $\sum_{i=1}^{M}Y_{i}=N$ is equivalent to the distribution of the balls-into-bins problem $(X_{1},\dots,X_{M})$ (see, e.g., Ref. [60]). Accordingly, if it is exponentially unlikely that the number of $Y_{i}=1$ is smaller than $cN^{\lambda}$ as we sample $(Y_{1},\dots,Y_{M})$, this indicates that it is still exponentially unlikely that the number of $X_{i}=1$ is smaller than $cN^{\lambda}$ as we sample $(X_{1},\dots,X_{M})$, given that the post-selection probability for $\sum_{i=1}^{M}Y_{i}=N$ is at most inverse polynomial (i.e., $\Theta(\frac{1}{\sqrt{N}})$).
@@
We emphasize that the distribution of the classical balls-into-bins problem is not equivalent to the uniform distribution over the possible outcomes of Boson Sampling, because the classical balls are distinguishable while bosons are indistinguishable in nature. Nevertheless, this example is sufficient to provide an insight that a sufficiently large number of singletons would appear in a large fraction of Boson Sampling outcomes.
Appendix H Numerical simulation for the collision patterns of Boson Sampling for the local random circuit over $(\mathcal{B}\mathcal{B}^{*})^{q}$
@@
Figure 6: The average ratio of collision-free outcomes from (a) $256$-mode and (b) $32$-mode unitary circuits drawn from local random circuit ensemble $\mathcal{H}_{\mathcal{A}}$ over the $q$-Kaleidoscope circuit architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ with the repetition number $q=1,2,3$, and global Haar random circuit ensemble. We sample 500 unitary circuits from each of the circuit ensembles, and we sample 500 outcomes on input $N=4,8,12,16$ photons for each of the randomly chosen circuits. Then we count the collision-free outcomes over the 500 sampled outcomes and compute the average ratio of collision-free outcomes over 500 randomly chosen circuits for each of the circuit ensembles. We present in the figure the average ratio of collision-free outcomes for local random circuits over $(\mathcal{B}\mathcal{B}^{*})^{q}$ with $q=1,2,3$ and global Haar random circuit ensemble, for the photon number given by $N=4,8,12,16$ (the error bar represents standard deviation). For all of the repetition numbers $q=1,2,3$ and the photon numbers $N=4,8,12,16$, local random circuits for $(\mathcal{B}\mathcal{B}^{*})^{q}$ exhibits almost equivalent average collision-free ratio to the global Haar random unitary circuits.
@@
To show the bosonic birthday paradox property for the local random circuit ensemble over $(\mathcal{B}\mathcal{B}^{*})^{q}$, we show numerically that the local random circuit ensemble over $(\mathcal{B}\mathcal{B}^{*})^{q}$ samples sufficiently many collision-free outcomes over the experiments. Specifically, we prepare random unitary circuits drawn from the local random circuit ensemble over $(\mathcal{B}\mathcal{B}^{*})^{q}$, sample the outcomes by classical Boson Sampling algorithm for each randomly chosen circuit, and compute the ratio of the collision-free outcomes over all the sampled outcomes for each randomly chosen circuit. To evaluate whether these local random circuits output sufficiently many collision-free outcomes, as a comparison, we also prepare the global Haar random circuits, which have the bosonic birthday paradox property as depicted in [1]. In other words, if the ratio of collision-free outcomes for local random circuit ensemble over $(\mathcal{B}\mathcal{B}^{*})^{q}$ is as large as the ratio of collision-free outcomes for global Haar random circuit, this clearly gives evidence of the bosonic birthday paradox for our circuit architecture.
@@
More specifically, we consider $256$-mode random unitary circuits drawn from local random circuit ensemble $\mathcal{H}_{\mathcal{A}}$ over the $q$-Kaleidoscope circuit architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ with the repetition number $q=1,2,3$, and global Haar random circuit ensemble. First, we sample 500 unitary circuits from each of the circuit ensembles described above. For each of the randomly chosen circuits, on input $N=4,8,12,16$ photons in the first $N$ modes over $256$ modes, we sample 500 outcomes using the Clifford-Clifford algorithm for Boson Sampling [61]. Then, we count the number of collision-free outcomes over the 500 sampled outcomes and compute the average ratio of collision-free outcomes for each $N$ and circuit ensemble, where the average is taken over the $500$ sampled unitary circuits for each circuit ensemble.
@@
In Fig. 6, we present the average ratio of collision-free outcomes and standard deviation (error bar) for the photon number given by $N=4,8,12,16$, from the local random circuits over $(\mathcal{B}\mathcal{B}^{*})^{q}$ with the repetition number $q=1,2,3$ and global Haar random circuits. Remarkably, local random circuits for $(\mathcal{B}\mathcal{B}^{*})^{q}$ shows great resemblance to the global Haar random circuits: For all of the photon numbers $N=4,8,12,16$ and the repetition numbers $q=1,2,3$, local random circuits for $(\mathcal{B}\mathcal{B}^{*})^{q}$ exhibits almost identical average collision-free ratio (and also the standard deviation) to the global Haar random circuits. Given that the bosonic birthday paradox is guaranteed for the global Haar random circuit ensemble, also given that the mode number we used covers all recent large-sized Boson Sampling experiments [4, 5, 6, 7], our numerical result shows that the local random circuit ensemble for $(\mathcal{B}\mathcal{B}^{*})^{q}$ guarantees the bosonic birthday paradox at least for the near-term finite-sized experiments. Even though this does not guarantee the bosonic birthday paradox of the local random circuit ensemble for $(\mathcal{B}\mathcal{B}^{*})^{q}$ in the asymptotic limit as the system size scales, still, we believe that constant repetition number $q$ would be sufficient for the bosonic birthday paradox, due to the close resemblance to the global Haar random circuit ensemble as shown in the figure.
@@
Additionally, we similarly present a numerical simulation result in Fig. 6 for the mode number $M=32$ and photon number $N=4,8,12,16$, which indicates the saturated regime where the bosonic birthday paradox does not hold. Even in this regime, the collision patterns of local random circuits over $(\mathcal{B}\mathcal{B}^{*})^{q}$ also show a great resemblance to those of global Haar random circuits. This result gives hope that one can possibly extend our result to the saturated regime, by conducting a similar analysis for our circuit ensemble to the analysis for global Haar random circuits [29].

### v3

@@ Appendix F Proof of Theorem 4 (average-case hardness)
For the polynomial interpolation, we employ the Robust Berlekamp-Welch algorithm that requires the NP oracle, recently proposed in Ref. [bouland2022noise].
Theorem 10 (Robust Berlekamp-Welch [bouland2022noise]).
Let $P(x)$ be a degree $d$ polynomial in $x$. Suppose there is a set of points $D=\{(x_{i},y_{i})\}$ such that $|D|=O(d^{2})$ and $\{x_{i}\}$ is equally spaced in the interval $0,\Delta]$. Suppose also that each points $(x_{i},y_{i})$ satisfies
@@
Appendix G Proof of Lemma 3
Let $\bar{p}_{\bm{s}}(C)$ be the output probability distribution from the approximate sampler $\mathcal{S}$ with the given linear optical circuit $C$. Also, let $\mathcal{C}_{M,N}$ be the set of collision-free outcomes of Boson Sampling, for mode number $M$ and photon number NN. Then $\bar{p}_{\bm{s}}(C)$ satisfies
@@ Appendix G Proof of Lemma 3
for all $k>2$. Also, using Stockmeyer’s algorithm [stockmeyer1985approximation] whose complexity is in $\rm{BPP}^{\rm{NP}}$, obtaining the estimate $\tilde{p}_{\bm{s}}(C)$ of $\bar{p}_{\bm{s}}(C)$ satisfying
@@
in polynomial time in NN and $\alpha^{-1}$ is in $\rm{BPP}^{\rm{NP}^{\mathcal{S}}}$. Using $\operatorname*{\mathbb{E}}_{\bm{s}\sim\mathcal{G}_{M,N}}[\bar{p}_{\bm{s}}(C)]=\binom{M}{N}^{-1}\sum_{\bm{s}\in\mathcal{C}_{M,N}}\bar{p}_{\bm{s}}(C)\leq\binom{M}{N}^{-1}$,
@@
Appendix H On average-case hardness including collision outcomes
@@
As we have remarked in the main text, extending our hardness result of shallow-depth Boson Sampling to the saturated regime $M=\Theta(N^{\gamma})$ for $1\leq\gamma<2$ remains a crucial open problem, and to do so, one also needs to consider collision outcomes in the average-case hardness argument. In this appendix, we present a possible approach that can be further improved to the average-case hardness result including collision outcomes.
@@
Our intuition is that while most of the outcomes in the saturated regime are not collision-free, the key point is that most outcomes are expected to be “nearly" collision-free, in the sense that for most of the outcomes, a large portion of the measured modes are occupied only by a single photon. Also, we can establish worst-to-average-case reduction across different system sizes (i.e., mode number & photon number), i.e., the system size for the worst-case hardness result can be smaller at most polynomially than the system size for the average-case hardness result. Hence, by post-selecting outcomes with the number of measured single photons at least as large as the photon number required in the worst-case problem (in Theorem 3, which requires a collision-free outcome), one would possibly establish worst-to-average-case reduction from the randomly chosen outcome over all possible outcomes.
@@
More specifically, let the mode number $M=\Theta(N^{\gamma})$ for any $\gamma\geq 1$. Let us define $\mathcal{Q}_{M,N}$ as the uniform distribution over all the possible outcomes of Boson Sampling with $M$ modes and NN photons, such that each outcome $\bm{s}\sim\mathcal{Q}_{M,N}$ is an $M$-dimensional output configuration vector $\bm{s}=(s_{1},\dots,s_{M})$ with the constraint $\sum_{i=1}^{M}s_{i}=N$. Also, we denote the number of $s_{i}=1$ in $\bm{s}$ as the number of singletons throughout this appendix. Here, let us assume that there exists $N_{0}=cN^{\lambda}$ for a constant $c$ and $0<\lambda<1$ such that it is likely that the number of singletons is more than $N_{0}$; we later provide the intuition to justify this assumption at the end of this appendix.
@@
Hereafter, we denote each $(M_{0},N_{0})$ and $(M,N)$ as the system sizes (mode number & photon number) for the worst-case hardness (i.e., Theorem 3) and average-case hardness (we aim to prove), respectively. More specifically, given a worst-case circuit $C_{0}$ in $M_{0}=\Omega(N_{0})$ mode circuit architecture $\mathcal{B}\mathcal{B}^{*}$ and a fixed $M_{0}$-mode and $N_{0}$-photon collision-free outcome $\bm{s}_{0}$ depicted in Theorem 3, by our assumption, one can find a photon number NN that satisfies $N_{0}=cN^{\lambda}$ for a constant $c$ and $0<\lambda<1$, and mode number $M=\Theta(N^{\gamma})$ for $\gamma\geq 1$, such that most of the outcomes over $\mathcal{Q}_{M,N}$ have at least $N_{0}$ singletons. In the following, we show a possible approach for worst-to-average-case reduction for general outcomes, by describing how one can estimate $p_{\bm{s}_{0}}(C_{0})$ to within desired imprecision given oracle access to the well-estimated values $p_{\bm{s}}(U)$ for average-case outcomes $\bm{s}\sim\mathcal{Q}_{M,N}$ and $M$-mode average-case circuits over $U\sim\mathcal{H}_{\mathcal{A}}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$.
@@
We now argue how to estimate the worst-case value $p_{\bm{s}}(C)$ and the average-case value $p_{\bm{s}_{1}}(U_{0})$. First, estimating the average-case value $p_{\bm{s}_{1}}(U_{0})$ can be done by similarly accessing the oracle, but now we input the outcome $\bm{s}_{1}$ and the circuit $U_{0}$ instead. Hence, the success probability for the oracle would increase at most the total variation distance between the distribution of $\bm{s}_{1}$ and $\mathcal{Q}_{M/2,N-N_{0}}$, and if this can be sufficiently bounded, we can obtain the estimated value of $p_{\bm{s}_{1}}(U_{0})$ with high probability. Also, one can estimate the value $p_{\bm{s}}(C)$ by the polynomial interpolation, as in the proof of Theorem 4. Specifically, by parameterizing the circuit $U$ as $U(\theta)$, where each gate of $U$ (except those in $U_{0}$) is perturbed via the Cayley transform using the corresponding gate from $C$, one can similarly infer the value $p_{\bm{s}}(C)$ through oracle access to $p_{\bm{s}}(U(\theta))$ and applying polynomial interpolation (see Appendix F for details).
@@
We highlight the necessary ingredients that need to be addressed for the successful worst-to-average-case reduction by the above process. First, to make sure that we can also well-estimate $p_{\bm{s}_{1}}(U_{0})$ with high probability over $\bm{s}_{1}$ via the oracle access, one needs to justify that the distribution of $\bm{s}_{1}$ is sufficiently close to the distribution $\mathcal{Q}_{M/2,N-N_{0}}$. To help understand how the distribution of $\bm{s}_{1}$ can be close to $\mathcal{Q}_{M/2,N-N_{0}}$, consider the limiting case that $N_{0}=0$. Here, given that $M$ is much larger than NN (but still can be $M=\Theta(N)$), sampling $\bm{s}_{1}$ is equivalent to sampling from $\mathcal{Q}_{M,N}$ and post-selecting outcomes that have all the non-zero elements in the first $M/2$ mode, which yields the uniform distribution $\mathcal{Q}_{M/2,N}$. Therefore, given that $M$ is much larger than NN and $N_{0}$ is sufficiently small relative to NN, we expect that the distribution of $\bm{s}_{1}$ would not largely deviate from the distribution $\mathcal{Q}_{M/2,N-N_{0}}$.
@@
Also, as the final imprecision level increases by the estimated value of $p_{\bm{s}_{1}}(U_{0})^{-1}$, it must be promised that $p_{\bm{s}_{1}}(U_{0})$ is lower bounded. More precisely, $p_{\bm{s}_{1}}(U_{0})$ must be at least as large as the allowed imprecision level of average-case estimation (i.e., imprecision level of the oracle $\mathcal{O}$), commonly referred to as the anti-concentration property [aaronson2011computational, bouland2019complexity, Bremner2016average, Hangleiter2018anticoncentration, Bermejo2018architectures, Dalzell2022random]. Nevertheless, we believe that $p_{\bm{s}_{1}}(U_{0})$ is at least as large as the allowed imprecision level for most of $U_{0}$, based on two reasons: First, the imprecision level required by our current average-case hardness result is sufficiently small. Also, although this has only been demonstrated in finite-sized numerical experiments, random circuits for our circuit architecture closely resemble global Haar-random unitary circuits [go2024exploring], which are strongly conjectured to exhibit the anti-concentration property [aaronson2011computational].
@@
We now present a toy model that could offer an insight that, for a large portion of outcomes of Boson Sampling, a sufficient number of singletons would exist. In the following, we present the classical “balls-into-bins" problem where we throw NN balls into $M$ bins uniformly at random, which is indeed the classical analogue of Boson Sampling. We show that, given that $M=\Theta(N^{\gamma})$ with $\gamma\geq 1$, it is highly unlikely that the number of bins getting exactly one ball is less than $cN^{\lambda}$ for a constant $c$ and $0<\lambda<1$.
@@
Let us first denote the random variable $X_{i}$ as the number of balls in $i$-th bin (such that $\sum_{i=1}^{M}X_{i}=N$), when we throw NN balls into $M$ bins uniformly at random. Here, our goal is to verify that over the experiments, it is highly unlikely to observe the outcomes that the number of $X_{i}=1$ is smaller than $cN^{\lambda}$. Since all the variables $X_{i}$ are correlated due to the constraint $\sum_{i=1}^{M}X_{i}=N$, instead, we consider i.i.d. random variable $Y_{i}$ each drawn from Poisson distribution $P(\frac{N}{M})$, such that $\operatorname*{\mathbb{E}}[\sum_{i=1}^{M}Y_{i}]=N$ and $\Pr[\sum_{i=1}^{M}Y_{i}=N]=e^{-N}\frac{N^{N}}{N!}=\Theta(\frac{1}{\sqrt{N}})$. The motivation for using Poissonian random variables is that the distribution of $(Y_{1},\dots,Y_{M})$ conditioned on the sum $\sum_{i=1}^{M}Y_{i}=N$ is equivalent to the distribution of the balls-into-bins problem $(X_{1},\dots,X_{M})$ (see, e.g., Ref. [ross2014introduction]). Accordingly, if it is exponentially unlikely that the number of $Y_{i}=1$ is smaller than $cN^{\lambda}$ as we sample $(Y_{1},\dots,Y_{M})$, this indicates that it is still exponentially unlikely that the number of $X_{i}=1$ is smaller than $cN^{\lambda}$ as we sample $(X_{1},\dots,X_{M})$, given that the post-selection probability for $\sum_{i=1}^{M}Y_{i}=N$ is at most inverse polynomial (i.e., $\Theta(\frac{1}{\sqrt{N}})$).
@@
We emphasize that the distribution of the classical balls-into-bins problem is not equivalent to the uniform distribution over the possible outcomes of Boson Sampling, because the classical balls are distinguishable while identical bosons are indistinguishable in nature (and indeed, bosons are even more “gregarious" than classical balls [aaronson2011computational]). Nevertheless, this example is sufficient to provide an insight that a sufficiently large number of singletons would appear in a large fraction of Boson Sampling outcomes.
We conclude this appendix by introducing an alternative approach, based on the closely related result in [bouland2023complexity] that establishes a hardness argument for general Boson Sampling in the saturated regime. Note that our convention on singletons closely parallels the notion of “clicks" introduced in [bouland2023complexity], which refers to the number of modes that are occupied by a non-zero number of photons. Analogous to our analysis, their key observation is that, even in the saturated regime, most outcomes have a sufficiently large number of clicks. Based on this observation, they show that, given an NN-photon outcome $\bm{s}$ with collisions that has a sufficient number of clicks $N_{0}$, there exists a poly-time constructible matrix $A\in\{0,1\}^{N_{0}\times N}$ such that computing $\text{Per}(A_{\bm{s}})$ is worst-case #P-hard, for $A_{\bm{s}}$ being a submatrix of AA by repeating rows according to $\bm{s}$. Then they establish a worst-to-average-case reduction by perturbing AA into their average-case matrix and performing polynomial interpolation (for details, see [bouland2023complexity]).
Accordingly, since we employ a conceptually similar strategy in the main text, establishing worst-case hardness for shallow-depth circuits with a large enough number of clicks is sufficient to obtain an analogous hardness result for shallow-depth Boson Sampling. This requires constructing a shallow-depth circuit (as done in the proof of Theorem 3) for which estimating an output probability with a large enough number of clicks is worst-case #P-hard, in direct analogy to the construction of the worst-case matrix AA above. Since this lies beyond the scope of the presented work, we leave this as future work.
Appendix I Numerical simulation for the collision patterns of Boson Sampling for the local random circuit over $(\mathcal{B}\mathcal{B}^{*})^{q}$
@@
To show the bosonic birthday paradox property for the local random circuit ensemble over $(\mathcal{B}\mathcal{B}^{*})^{q}$, we show numerically that the local random circuit ensemble over $(\mathcal{B}\mathcal{B}^{*})^{q}$ samples sufficiently many collision-free outcomes over the experiments. Specifically, we prepare random unitary circuits drawn from the local random circuit ensemble over $(\mathcal{B}\mathcal{B}^{*})^{q}$, sample the outcomes by classical Boson Sampling algorithm for each randomly chosen circuit, and compute the ratio of the collision-free outcomes over all the sampled outcomes for each randomly chosen circuit. To evaluate whether these local random circuits output sufficiently many collision-free outcomes, as a comparison, we also prepare the global Haar random circuits, which have the bosonic birthday paradox property as depicted in [aaronson2011computational]. In other words, if the ratio of collision-free outcomes for local random circuit ensemble over $(\mathcal{B}\mathcal{B}^{*})^{q}$ is as large as the ratio of collision-free outcomes for global Haar random circuit, this clearly gives evidence of the bosonic birthday paradox for our circuit architecture.
@@
Figure 6: The average ratio of collision-free outcomes from (a) 256256-mode and (b) 3232-mode unitary circuits drawn from local random circuit ensemble $\mathcal{H}_{\mathcal{A}}$ over the $q$-Kaleidoscope circuit architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ with the repetition number $q=1,2,3$, and global Haar random circuit ensemble. We sample 500 unitary circuits from each of the circuit ensembles, and we sample 500 outcomes on input $N=4,8,12,16$ photons for each of the randomly chosen circuits. Then we count the collision-free outcomes over the 500 sampled outcomes and compute the average ratio of collision-free outcomes over 500 randomly chosen circuits for each of the circuit ensembles. We present in the figure the average ratio of collision-free outcomes for local random circuits over $(\mathcal{B}\mathcal{B}^{*})^{q}$ with $q=1,2,3$ and global Haar random circuit ensemble, for the photon number given by $N=4,8,12,16$ (the error bar represents standard deviation). For all of the repetition numbers $q=1,2,3$ and the photon numbers $N=4,8,12,16$, local random circuits for $(\mathcal{B}\mathcal{B}^{*})^{q}$ exhibits almost equivalent average collision-free ratio to the global Haar random unitary circuits.
More specifically, we consider 256256-mode random unitary circuits drawn from local random circuit ensemble $\mathcal{H}_{\mathcal{A}}$ over the $q$-Kaleidoscope circuit architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ with the repetition number $q=1,2,3$, and global Haar random circuit ensemble. First, we sample 500 unitary circuits from each of the circuit ensembles described above. For each of the randomly chosen circuits, on input $N=4,8,12,16$ photons in the first NN modes over 256256 modes, we sample 500 outcomes using the Clifford-Clifford algorithm for Boson Sampling [clifford2018classical]. Then, we count the number of collision-free outcomes over the 500 sampled outcomes and compute the average ratio of collision-free outcomes for each NN and circuit ensemble, where the average is taken over the 500500 sampled unitary circuits for each circuit ensemble.
@@
In Fig. 6, we present the average ratio of collision-free outcomes and standard deviation (error bar) for the photon number given by $N=4,8,12,16$, from the local random circuits over $(\mathcal{B}\mathcal{B}^{*})^{q}$ with the repetition number $q=1,2,3$ and global Haar random circuits. Remarkably, local random circuits for $(\mathcal{B}\mathcal{B}^{*})^{q}$ shows great resemblance to the global Haar random circuits: For all of the photon numbers $N=4,8,12,16$ and the repetition numbers $q=1,2,3$, local random circuits for $(\mathcal{B}\mathcal{B}^{*})^{q}$ exhibits almost identical average collision-free ratio (and also the standard deviation) to the global Haar random circuits. Given that the bosonic birthday paradox is guaranteed for the global Haar random circuit ensemble, also given that the mode number we used covers all recent large-sized Boson Sampling experiments [zhong2020quantum, zhong2021phase, madsen2022quantum, deng2023gaussian], our numerical result shows that the local random circuit ensemble for $(\mathcal{B}\mathcal{B}^{*})^{q}$ guarantees the bosonic birthday paradox at least for the near-term finite-sized experiments. Even though this does not guarantee the bosonic birthday paradox of the local random circuit ensemble for $(\mathcal{B}\mathcal{B}^{*})^{q}$ in the asymptotic limit as the system size scales, still, we believe that constant repetition number $q$ would be sufficient for the bosonic birthday paradox, due to the close resemblance to the global Haar random circuit ensemble as shown in the figure.
@@
Additionally, we similarly present a numerical simulation result in Fig. 6 for the mode number $M=32$ and photon number $N=4,8,12,16$, which indicates the saturated regime where the bosonic birthday paradox does not hold. Even in this regime, the collision patterns of local random circuits over $(\mathcal{B}\mathcal{B}^{*})^{q}$ also show a great resemblance to those of global Haar random circuits. This result gives hope that one can possibly extend our result to the saturated regime, by conducting a similar analysis for our circuit ensemble to the analysis for global Haar random circuits [bouland2023complexity].